# On-Orbit Servicing-Integrated Maintenance Strategy for Satellite Constellation

### v1

@@ Abstract
This paper proposes a maintenance strategy for a satellite constellation that utilizes on-orbit servicing (OOS). Under this strategy, the constellation operator addresses satellite failures in two ways: by deploying new satellites and by recovering failed satellites through OOS. We develop an inventory management model with a parametric replenishment policy for the maintenance process, which can evaluate the performance of the satellite constellation system. Based on this model, we formulate two single-objective optimization problems representing the decision-making contexts of two main stakeholders—the constellation operator and the OOS provider—and a bi-objective optimization problem that can reflect the tension between the two. A case study of the OOS-supported maintenance for a real-world-scale constellation provides valuable insights that help explain the behaviors of the stakeholders.
@@
$\mathbf{x}_{\mathrm{con}}$  = decision variables vector of constellation operator
$\mathbf{x}_{\mathrm{oos}}$  = decision variables vector of OOS
@@ 1 Introduction
Satellite mega-constellations have played a pivotal role in shaping the modern space industry, driven by the demand for global coverage services and advances in space technologies. These driving forces have encouraged active participation from major players, including SpaceX’s Starlink, Amazon’s Project Kuiper, and Eutelsat’s OneWeb [1, 2]. This participation has expanded the market and established mega-constellations as a critical component in the ecosystem of the space industry.
From a system maintenance perspective, satellite mega-constellations expose operators to more frequent component failures (i.e., satellite failures), primarily due to the reduced reliability of individual satellites and the vastly increased number of units in operation. Addressing these challenges necessitates a systematic maintenance strategy, specifically, a well-structured logistics framework for effective failure response [3]. Recent studies, including those by Jakob et al. [4] and Kim et al. [5], have proposed strategies and corresponding assessment tools to tackle this problem.
@@
Beyond deploying spare satellites, on-orbit servicing (OOS) can recover a satellite’s operational capability and is beginning to be commercialized [8, 9, 10, 11]. Historically, OOS has targeted heavy space systems, such as the International Space Station, the Hubble Space Telescope, and communication satellites at geosynchronous equatorial orbit, whose life extension offers significant monetary value. Numerous studies have quantified OOS’s benefits for single large systems [12, 13, 14, 15, 16, 17, 18], and operational successes, such as the MEV-1 mission [19], underscore its practical value.
More recently, researchers have begun to consider OOS for low Earth orbit (LEO) constellations, where the rapid growth of space activities presents new opportunities. Although each individual satellite is relatively inexpensive and the added value of servicing a single unit is smaller compared to larger space systems, the overall scale of constellations continues to attract players in the space industry. Luu and Hastings [20] compared two strategies—deploying spare satellites and using OOS—for maintaining satellite mega-constellations. However, existing research has not yet examined spare satellites and OOS simultaneously as part of a unified maintenance strategy for satellite mega-constellations.
@@
To address this gap, we propose a novel maintenance strategy for satellite constellations that integrates OOS with the deployment of spare satellites. Within this strategy, failed satellites are not only replaced by spare satellites but can also be repaired through OOS if the failure is recoverable. To evaluate the performance of the proposed strategy at a strategic level, we construct the corresponding supply chain and develop an assessment tool based on an inventory management model. Building on this foundation, we introduce relevant decision-making scenarios, formulate them as optimization problems, and conduct a case study to demonstrate the applicability of the proposed approach and provide actionable insights.
The remainder of this paper is organized into five sections. Section 2 introduces the modeling assumptions and constructs the supply chain and inventory management framework for evaluating the OOS-integrated maintenance strategy. Section 3 describes the decision-making scenarios, while Section 4 presents a case study that illustrates the application of these scenarios. Finally, Section 5 concludes the paper.
@@ 2 Model Construction
This section outlines the proposed maintenance strategy and the evaluation framework. We begin by introducing the supply chain and the underlying assumptions. Next, we present the parametric models that describe the mean behavior of the inventory at each echelon of the supply chain and construct the maintenance cost and system performance parameters. Finally, we assess the model’s accuracy by comparing its results with those obtained from discrete-time simulations.
@@ 2.1 Supply Chain Overview and Assumptions
We assume that the target system for maintenance has a Walker Delta pattern [21], with inclination $i$ and altitude $h_{\mathrm{plane}}$. It consists of $N_{\mathrm{plane}}$ orbital planes whose right ascensions of ascending nodes (RAANs) are equally spaced by $2\pi/N_{\mathrm{plane}}$. Each orbital plane contains $N_{\mathrm{sat}}$ operational satellites.
The operator sources spare satellites in two ways: 1) recovering the operational life of failed satellites via OOS, and 2) launching newly manufactured satellites from the ground. Figure 1 presents an overview of the proposed supply chain.
@@ 2.1.1 Sourcing Method 1: Life Extension by OOS
The service lead time, defined as the duration between service request reception and service completion, is primarily composed of a series of maneuvers performed by the servicing spacecraft to approach the target satellite (rendezvous, proximity operations, docking), along with the time required to execute the servicing activity itself. The lead time depends on the current state of the OOS provision system. For example, if an available servicing spacecraft is spatially distant from the target satellite at the time of the service request, the lead time will be long. Since the time required for the available servicing spacecraft to maneuver to the target is inherently random, the lead time is modeled as a probabilistic variable.
@@ 2.1.2 Sourcing Method 2: Deploying New Satellites
In-plane spares directly replace failed satellites and are distributed across the (near-)target orbits (i.e., the operational planes of the constellation). When the inventory of in-plane spares reaches a reorder point $s$, a batch of $Q$ spares is sourced from the parking spares, the next echelon. Here, the spare satellites are transported by leveraging $J_{2}$ perturbation to align RAANs of the parking and target orbits, and the satellite’s propulsion system to raise the altitude. Here, we have the RAAN drift rate relation [22]:
@@
    In-orbit failures of operational satellites follow a Poisson process with constant rate $\lambda_{\mathrm{sat}}$.
@@
    Spare satellites’ failures are neglected (i.e., ground, parking, and in‐plane spares are assumed failure-free).
@@
    In-plane spares are available instantaneously to replace a failed satellite [4].
@@
    Each target orbit is replenished from the closest available parking orbit (i.e., a nonempty parking orbit with the earliest RAAN alignment). When an in-plane stock level hits its reorder point ($s$), a batch of size $Q$ is drawn from that parking orbit [4].
@@
    Ground-to-parking delivery can supply only a single parking orbit [4].
@@
    A new in-plane order is placed only after the previous one has completed transit to facilitate the tracking of the orders [4].
@@
    Stock-outs are rare; therefore, the probability of all parking orbits being out of stock is negligible [4].
@@
    Both the reorder point and order quantity at parking orbits are constrained to integer multiples of the batch size ($Q$) [4].
@@
    An OOS lead time follows an exponential distribution, which is a stochastic distribution characterized by a single parameter, with mean time to recovery (or equivalently, mean time to repair, MTTR), $\mu_{\mathrm{oos}}^{-1}$.
@@ 2.2 In-Plane Spares
In-plane spares are replenished by parking spares and OOS. Figure 2 provides a closer view of the closed-loop system shown in Fig. 1, illustrating the flows of spare and failed satellites for the case of $N_{\mathrm{oos}}=2$. Since our model represents the mean behavior over an infinite time horizon, we define $\gamma_{0}$ as the fraction of newly arrived in-plane spares sourced from parking orbits, and $\gamma_{m}$ ($m=1,2,\dots,N_{\mathrm{oos}}$) as the fraction of in-plane spares that have been serviced $m$ times, also in the infinite time horizon.
@@
Assuming that the service lead time follows an exponential distribution with the mean value of $\mu_{\mathrm{oos}}^{-1}$ implies that the OOS process recovers a failed satellite with probability $\mu_{\mathrm{oos}}$ per unit time [25]. Based on this assumption, the flow balance equations at the steady-state for failed satellites awaiting OOS and in-plane spares over an infinite horizon are as follows:
@@
where $P_{\mathrm{av}}^{\prime}\left(j\right)$ is the normalized probability that the $j$th closest parking orbit is the closest available parking one, $\dot{\Omega}_{\mathrm{rel}}$ is the relative RAAN drift rate between target and parking orbits, and $\mathcal{T}_{j}$ is the admissible range of alignment time between a target orbit and its $j$th closest parking orbit, defined as:
@@
Here, $t_{\mathrm{trans}}$ is the time of flight from a parking orbit to a target orbit calculated by Eq. (3).

### v2

@@ Abstract
This paper proposes a maintenance strategy for a satellite constellation that utilizes on-orbit servicing (OOS). Under this strategy, the constellation operator addresses satellite failures in two ways: by deploying new satellites and by recovering failed satellites through OOS. We develop an inventory management model with a parametric replenishment policy for the maintenance process, which can evaluate the performance of the satellite constellation system. Based on this model, we formulate the interaction between the constellation operator—who seeks to maintain the required service level of the constellation while minimizing maintenance cost—and the OOS provider—who seeks to maximize profit by selecting service price and performance levels—as a bi-objective optimization problem and identify the corresponding Pareto-optimal solutions. A case study based on real-world-scale constellation and launch service shows that, relative to the benchmark strategy without OOS, the OOS-integrated solutions can reduce annual maintenance cost by up to 14.5%, while reducing annual launch and manufacturing costs by approximately 25% each and maintaining the required service levels. The results further show that, within a given scenario, the Pareto-optimal set is generally generated by an almost invariant maintenance strategy on the constellation operator side, whereas most variation along the Pareto frontier is driven by pricing decisions of the OOS provider side. Among the OOS-related parameters, the fraction of failures that can be recovered through OOS has the strongest structural effect.
@@
$\mathbf{x}$  = decision variables vector of optimization problem
@@ 1 Introduction
Satellite mega-constellations have played a pivotal role in shaping the modern space industry, driven by the demand for global coverage services and advances in space technologies. These driving forces have encouraged active participation from major players, including SpaceX’s Starlink, Amazon Leo, and Eutelsat’s OneWeb [1, 2]. This participation has expanded the market and established mega-constellations as a critical component in the ecosystem of the space industry.
From an operational perspective, mega-constellations in low Earth orbit (LEO) pose challenges throughout the lifecycle that differ substantially from those of traditional constellations in geostationary equatorial orbit (GEO). The large number of satellites, together with the relatively small size of individual spacecraft and tight constraints on power, performance, and maneuvering capability, complicates system operation and management. Prior research has addressed multiple lifecycle phases, including pattern design [3, 4, 5, 6, 7, 8, 9, 10, 11, 13, 12, 14, 15], deployment [16, 17, 18, 19, 20, 21], mission operations [22, 23, 24, 25], and end-of-life management [26, 27]. This study focuses on maintenance. In mega-constellations, component-level failures (i.e., in-orbit failures of individual satellites) may occur more frequently, not only because more satellites are in operation, but also because individual satellites are often designed under tighter mass, cost, and redundancy constraints. In addition, rapid technology refresh cycles may require agile integration of payload and spacecraft upgrades. For example, Starlink satellites are reported to have an operational lifetime of about five years, a wet mass of several hundred kilograms, and a planned constellation size of 42,000 [28]. In contrast, current Intelsat—a GEO-based telecommunication company—satellites can exceed 6 metric tons, are expected to remain operational for more than 15 years [29], and are deployed in fleets of only a few tens of satellites. These differences motivate the development of a systematic maintenance strategy tailored to mega-constellations, specifically a structured logistics framework for effective failure response [30]. Recent studies, including Jakob et al. [31] and Kim et al. [32], have proposed maintenance strategies and corresponding assessment tools for satellite mega-constellations.
@@
Beyond replacing failed satellites with new spare units, on-orbit servicing (OOS) can restore operational capability and is beginning to be commercialized [34, 35, 36, 37]. Historically, OOS has focused on high-value space systems such as the International Space Station, the Hubble Space Telescope, and communication satellites in GEO, for which life extension can provide substantial economic value. Numerous studies have assessed the benefits of OOS for single large systems [38, 39, 40, 41, 42, 43, 44], and operational missions such as MEV-1 [45] indicate its practical relevance.
More recently, studies on OOS for servicing multiple satellites have examined technology and architecture alternatives [50, 46, 51, 47, 52, 48, 49], as well as the optimization of service planning and system design [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. However, this body of literature primarily adopts the perspective of the OOS provider, focusing on how OOS systems—an emerging and still difficult-to-realize concept—should be designed and operated. By contrast, the customer-side perspective, namely how constellation operators should incorporate OOS into their system, remains relatively underexplored. For example, Luu and Hastings [52] compared two maintenance strategies for satellite mega-constellations from the constellation operator’s perspective: deploying spare satellites or relying on OOS. Nevertheless, to our knowledge, prior research has not jointly considered spare satellites and OOS within a unified maintenance strategy for satellite mega-constellations.
@@
To address this gap, this paper proposes a maintenance strategy for satellite constellations that integrates OOS with spare satellite management. Under this strategy, failed satellites may be replaced by spare units or recovered through OOS when the failure is recoverable. To evaluate the proposed strategy, we develop an assessment tool based on an inventory management model. Building on this model, we define a decision-making scenario involving two stakeholders: the constellation operator, as the customer of OOS, and the OOS provider. We then formulate these settings as optimization problems and conduct a case study.
The contributions of this work are twofold. First, we propose an inventory control-based framework for integrating OOS into mega-constellation maintenance systems. The framework jointly considers spare-based replacement and recovery through OOS, and incorporates the resulting closed-loop structure into the inventory management model. To our knowledge, this closed-loop feature has not been addressed in prior work [31, 32]. Second, building on this framework, we formulate an optimization problem and present a case study to demonstrate its applicability and derive managerial insights for relevant decision makers. In particular, we conduct a parametric analysis of OOS cost and performance, both of which remain uncertain as OOS capabilities continue to emerge. From the customer perspective (i.e., constellation operators), the analysis guides how operators should respond to the emergence of OOS in the market. From the customer perspective (i.e., constellation operators), the analysis clarifies how operators should respond to the emergence of OOS in the market. From the provider perspective (i.e., OOS providers), it identifies the key factors that make OOS offerings attractive to customers and examines how customers may respond to such offerings.
The remainder of this paper is organized as follows. Section 2 outlines the modeling assumptions and develops the supply chain and inventory management framework to evaluate the OOS-integrated maintenance strategy. Section 3 validates the proposed model by comparing its results with those obtained from inventory simulations and assesses its robustness to key assumptions through analyses under relaxed assumptions. Section 4 presents the decision-making scenarios and corresponding optimization problem formulations, and Section 5 provides a case study illustrating their application. Finally, Section 6 concludes the paper.
@@ 2 Model Construction
This section presents the proposed maintenance strategy and evaluation framework. We first describe the concept of operations of the maintenance system, together with the associated supply chain and its underlying assumptions. We then present the parametric models to characterize the mean inventory behavior at each echelon of the supply chain and relate these models to the maintenance cost and other system performance parameters.
2.1 Concept of Operations
Before presenting the detailed assumptions and inventory models, we briefly describe the maintenance system. Figure 1 provides an overview of the proposed maintenance concept.
Figure 1: Illustration of the proposed OOS-integrated maintenance strategy.
We assume that the target constellation follows a Walker Delta pattern [67], where all orbital planes are circular, share the same inclination, and are uniformly distributed over the right ascensions of the ascending nodes (RAANs). Satellites within each orbital plane are also assumed to be uniformly distributed along, and all satellites in the constellation are assumed to be homogeneous. Although the formulation is developed for this symmetric configuration, it can be adapted with minimal effort to constellations that deviate from it, for example, by allowing nonzero eccentricity in each orbital plane or nonuniform distribution of orbital planes over the RAAN domain. In such cases, the model can still be used to track the mean behavior of the spare inventory, although deviations from the mean would require additional study.
To address failures of operational satellites, we consider three echelons of spare satellites, following Jakob et al. [31]. The first echelon consists of in-plane spares, which are distributed across the (near-)target orbits (i.e., the operational planes of the constellation). When an operational satellite fails, its slot is filled by an in-plane spare. The second echelon consists of parking spares, which replenish the in-plane spares. These parking orbits have the same inclination as the target orbits and are uniformly distributed over RAANs, but they are located at lower altitudes. A parking spare replenishes an in-plane spare by performing an orbit-raising maneuver once the corresponding parking and target orbits become aligned in RAAN due to the differential RAAN drift induced by the altitude difference, as discussed later. The third echelon consists of ground spares, which replenish the parking spares. Ground spares are newly manufactured satellites that are delivered to the parking orbits by launch vehicles. The associated replenishment flows of newly manufactured satellites are shown as blue arrows in Fig. 1.
In addition to the replenishment flow of newly manufactured satellites, the maintenance system includes an OOS-based recovery process, represented by the red arrows in Fig. 1. If an operational satellite experiences a failure mode that is recoverable through OOS, the failed satellite remains near the target orbit of its original slot and waits for service. A servicer spacecraft deployed in advance by the OOS provider then performs OOS, after which the failed satellite is retained as an in-plane spare.
@@ 2.2 Supply Chain Overview and Assumptions
The satellite constellation considered in the proposed strategy is assumed to follow a Walker Delta pattern with inclination $i$ and altitude $h_{\mathrm{plane}}$. The constellation consists of $N_{\mathrm{plane}}$ orbital planes, whose RAANs are equally spaced by $2\pi/N_{\mathrm{plane}}$. Each orbital plane contains $N_{\mathrm{sat}}$ operational satellites.
As noted earlier, the constellation operator sources spare satellites through two channels: (1) recovering the operational life of failed satellites via OOS and (2) launching newly manufactured satellites from the ground. Figure 2 provides an overview of the proposed supply chain.
@@ 2.2.1 Sourcing Method 1: Life Extension by OOS
In this study, we consider the aspects most relevant to the customer, namely, service price and responsiveness. Specifically, we assume that the OOS provider has already deployed the infrastructure required to support the constellation under consideration, including orbital depots for fuel, replacement units, and other servicing resources, as well as multiple servicing spacecraft capable of approaching and restoring failed satellites. Under this assumption, the OOS system has sufficient capacity to respond to service requests with stable cost and performance.
The OOS price is the amount the constellation operator pays to use the OOS. In this study, OOS is charged per servicing event at a fixed price of $p_{\mathrm{oos}}$. Correspondingly, the OOS provider receives revenue of $p_{\mathrm{oos}}$ per servicing event. To provide OOS, the provider must allocate resources to each servicing event, and we model the corresponding per-event cost as $c_{\mathrm{oos}}$. As a result, the OOS provider’s profit per servicing event is $\left(p_{\mathrm{oos}}-c_{\mathrm{oos}}\right)$.
Service lead time is the representative measure of service responsiveness. It is defined as the duration between receipt of a service request and completion of the requested service, and it is the key performance parameter considered in this study. It consists primarily of the time required for the sequence of maneuvers by which a servicing spacecraft approaches the target satellite, including rendezvous, proximity operations, and docking, together with the time required to perform the servicing activity itself. The lead time depends on the current state of the OOS system. For example, if an available servicing spacecraft is far from the target satellite when the request is made, the resulting lead time is longer. Because the time required for a servicing spacecraft to reach the target is inherently uncertain, we model service lead time as a random variable. The modeling details are presented later in this section.
Note that the cost ($c_{\mathrm{oos}}$) and the responsiveness of OOS are closely related. Given the OOS technologies, service types, and target failure modes of the serviced spacecraft, the OOS provider’s operational effort determines both service responsiveness and cost. For example, if the OOS provider allocates more resources, such as additional servicing spacecraft, depots, and propellant for faster orbital transfers, the servicing cost increases, but so does responsiveness. This relationship is modeled in a simplified manner in this study and is introduced later in Section 4.
@@ 2.2.2 Sourcing Method 2: Deploying New Satellites
In-plane spares directly replace failed satellites. When the inventory of in-plane spares reaches a reorder point $s$, a batch of $Q$ spares is sourced from the parking spares, the next echelon. Here, the spare satellites are transported by leveraging $J_{2}$ perturbation to align RAANs of the parking and target orbits, and the satellite’s propulsion system to raise the altitude. Here, we have the RAAN drift rate relation [68]:
@@
    In-orbit failures of operational satellites follow a Poisson process with constant rate $\lambda_{\mathrm{sat}}$. This assumption is reasonable under the steady-state operation of mega-constellations, where individual satellite health variations are averaged out at the fleet level, yielding an approximately constant aggregate failure rate.
@@
    Failures of spare satellites are neglected; that is, parking and in-plane spares are assumed to be failure-free. This simplification is justified because the time spent in spare status is much shorter than the operational lifetime and can therefore be neglected at the strategic level.
@@
    In-plane spares are assumed to be available instantaneously to replace a failed satellite. This is reasonable because the associated response time is short, typically on the order of a few days [16], relative to other relevant time scales such as RAAN alignment, launch lead time, orbital transfer, and OOS lead time.
@@
    Each target orbit is replenished from the closest available parking orbit, defined as the nonempty parking orbit with the earliest RAAN alignment. When the inventory of in-plane spare reaches the reorder point, $s$, a batch of size $Q$ is dispatched from that parking orbit.
@@
    Ground-to-parking delivery is assumed to supply only a single parking orbit. Without an orbital transfer vehicle or additional mechanisms, supplying multiple parking orbits would be inefficient [71].
@@
    A new in-plane replenishment order is placed only after the previous order has completed transit to facilitate the tracking of the orders.
@@
    Stock-outs are assumed to be rare. Accordingly, the probability that all parking orbits are simultaneously out of stock is treated as negligible, consistent with nominal maintenance system operations.
@@
    The reorder point and order quantity at parking orbits are restricted to integer multiples of the batch size ($Q$).
@@
    The OOS system is assumed to have sufficient capacity to serve the target constellation, so that individual service requests may be treated independently and significant waiting time is rare. Under this assumption, OOS lead time is modeled as an exponential random variable. This choice is motivated by its common use for lead times with these characteristics and by its mathematical tractability [33, 72]. The corresponding mean time to recovery, or equivalently mean time to repair (MTTR), is given by $\mu_{\mathrm{oos}}^{-1}$.
Assumptions 1–8 are adapted from prior works [31, 32]. These assumptions are appropriate for strategic-level analysis of mega-constellations because they preserve the key operational characteristics while reducing the model’s complexity. Assumption 9 is newly introduced in this study. In the model validation (Section 3), we relax this assumption to examine whether the proposed model remains robust under alternative lead time conditions.
@@ 2.3 In-Plane Spares
In-plane spares are replenished by parking spares and OOS. Figure 3 provides a closer view of the closed-loop system shown in Fig. 2, illustrating the flows of spare and failed satellites for the case of $N_{\mathrm{oos}}=2$, meaning that a satellite can be recovered at most twice. From the perspective of replacing the slot vacated by a failed satellite, a spare newly transferred from the parking orbit and a spare recovered through OOS are operationally equivalent. However, for modeling purposes, we distinguish in-plane spares according to their service history, namely whether they are newly arrived from the parking orbit or have been recovered through OOS, and, in the latter case, by the number of times they have been serviced.
After a parking spare arrives in the target orbit, it remains as an in-plane spare until an operational satellite fails and vacates a slot. The spare then fills that slot and operates until it fails. If the failure is serviceable, which occurs for a fraction $r_{\mathrm{oos}}$ of failures, the failed satellite waits for OOS and returns as an in-plane spare after the service lead time. Otherwise, it is disposed of. This process repeats until the satellite experiences a non-serviceable failure or reaches the maximum allowable number of services, $N_{\mathrm{oos}}$.
Over an infinite horizon, let $\gamma_{0},\gamma_{1},\dots,\gamma_{N_{\mathrm{oos}}}$ denote the steady-state fractions of in-plane spares. Here, $\gamma_{0}$ denotes newly arrived spares from parking orbits, while $\gamma_{m}$ ($m=1,2,\dots,N_{\mathrm{oos}}$) denotes spares that have undergone service exactly $m$ times. Because in-plane spares wait to replace failed operational satellites, they are held as spare inventory. We denote their mean stock level by $\overline{\mathrm{SL}}_{\mathrm{plane}}$, which includes both newly arrived spares from parking orbits and spares that have been serviced any number of times. Likewise, serviceable failed satellites are held while awaiting service. We denote the corresponding mean stock level for satellites awaiting the $m$th service by $\overline{\mathrm{SL}}_{\mathrm{wait},m}$.
@@
Assuming that the service lead time follows an exponential distribution with the mean value of $\mu_{\mathrm{oos}}^{-1}$ implies that the OOS process recovers a failed satellite with probability $\mu_{\mathrm{oos}}$ per unit time [72]. The corresponding steady-state flow balance equations over an infinite horizon for 1) failed satellites awaiting the $m$th OOS and 2) in-plane spares that have been serviced $m$ times are indicated in Fig. 3 by the star pairs labeled 1-2 and 2-3, respectively. Within each color group (yellow and green), adjacent star pairs represent individual flow-balance relations, corresponding to Eqs. (6) and (7):
@@
where $P_{\mathrm{av}}^{\prime}\left(j\right)$ is the normalized probability that the $j$th closest parking orbit is the closest available parking one, $\dot{\Omega}_{\mathrm{rel}}$ is the relative RAAN drift rate between target and parking orbits, and $\mathcal{T}_{j}$ denotes the admissible range of lead times for sourcing from the $j$th closest parking orbit, including both the alignment time between the target orbit and its $j$th closest parking orbit and the orbit-raising time from the parking orbit to the target orbit, defined as:
@@
Here, $t_{\mathrm{trans}}$ is the time of flight from a parking orbit to a target orbit calculated by Eq. (3), and the right-open interval of $\mathcal{T}_{j}$ is used to avoid overlap between adjacent admissible intervals, thereby preventing double counting at the boundary.